# Mad Libs Are All You Need: Augmenting Cross-Domain Document-Level Event Argument Data

### v1

@@ Abstract
Document-Level Event Argument Extraction (DocEAE) is an extremely difficult information extraction problem — with significant limitations in low-resource cross-domain settings. To address this problem, we introduce Mad Lib Aug (MLA), a novel generative DocEAE data augmentation framework. Our approach leverages the intuition that Mad Libs, which are categorically masked documents used as a part of a popular game, can be generated and solved by LLMs to produce data for DocEAE. Using MLA, we achieve a 2.6-point average improvement in overall F1 score. Moreover, this approach achieves a 3.9 and 5.2 point average increase in zero and few-shot event roles compared to augmentation-free baselines across all experiments.
To better facilitate analysis of cross-domain DocEAE, we additionally introduce a new metric, Role-Depth F1 (RDF1), which uses statistical depth to identify roles in the target domain which are semantic outliers with respect to roles observed in the source domain. Our experiments show that MLA augmentation can boost RDF1 performance by an average of 5.85 points compared to non-augmented datasets.
@@ 1 Introduction
In recent years, there has been considerable progress in the domain of Event Argument Extraction (EAE), as advances in question answering Du and Cardie (2020), prompt tuning Ma et al. (2022), and semantic graph modeling Yang et al. (2023b) have led to state-of-the-art results on EAE benchmarks. However, some simplifying assumptions made in prior work limit their applicability in a real-world setting, including use of sentence-level data, reliance on large-scale annotation, or that training and testing data come from the same domain and thus share a similar distribution.
@@
Figure 1: Mad Libs are a classic game where a humorous story is created by (i) masking phrases from a document (ii) assigning high-level categories to the masked phrases, and (iii) someone who has not read the document fills-in the masked phrases given only the category information. In this paper, we connect the concept of Mad Libs to Event Argument Extraction, showing how generating and solving Mad Libs using LLMs creates high-quality EAE data augmentation.
@@
A key issue in the FSCD DocEAE problem is the need to model 0-shot roles Yang et al. (2023a), i.e. roles which exist in an event’s ontology and thus might appear in the test set, but have no representation in the training data. To illustrate this, consider the DocEE dataset Tong et al. (2022), where each new event in the FSCD training set only has 5 documents available for training. Across all samples representing the Volcano Eruption event, for example, zero documents contain annotation for the role The State of the Volcano (Dormant or Active). This makes this role impossible for token classification models to predict, and requires span extraction models such as question answering to rely on pre-trained knowledge to predict such roles. Few-shot roles in FSCD are also extremely challenging to model and similarly rely on generalizing knowledge learned during pre-training.
@@
One solution to this problem is to generate augmented training samples which increase representation of zero and few-shot roles. However, to the best of our knowledge, no prior work has generated DocEAE training samples in a way which can help resolve the aforementioned issues. For example, many related works in EAE data augmentation such as Gao et al. (2022); Ma et al. (2023); Wang et al. (2023) only focus on sentence-level tasks. The few prior works in DocEAE augmentation do not generate new data samples but rather augment existing samples Liu et al. (2022) or use a pre-trained model to weakly label unannotated corpora Liu et al. (2021).
@@
In this paper, we address this issue by introducing Mad Lib Aug (MLA), a novel, LLM-powered data augmentation framework for document-level EAE. To the best of our knowledge, MLA is the first data augmentation framework for generating data for event-centric document-level argument extraction task. MLA is inspired by the insight that EAE samples can be formulated as a Mad Lib, where the “blanks" found in Mad Libs correspond to roles in an event ontology. We visualize the relationship between Mad Libs and EAE in Figure 1. As shown in Figure 2, we use LLMs to both generate Mad Libs, and then solve them by generating JSON objects, which map event roles to textual arguments. Our method is advantageous in that we (i) formulate EAE data generation in the context of a popular game, Mad Libs, thus leveraging LLMs strong conceptual priors, (ii) our method requires no in-domain data and can generate data for 0-shot roles given only the event title and role name thus facilitating cross-domain EAE, (iii) augmented roles are inserted into Mad Lib “blanks" using string replacement, which is well-suited to extractive tasks. This point is crucial for EAE document generation, as trying to assert specific role arguments occur (or do not occur) during autoregressive generation of a long document can be difficult to perform reliably.
Figure 2: Visualization of our data generation framework. In the Document Generation stage, the LLM treats data generation as a Mad Lib generation task, with instructions to use Mad Lib categories which correspond to event roles. In the role-argument generation stage, the model is instructed to generate a JSON file with the missing information in each (category) slot in the synthesized Mad Lib. An Event Argument Extraction sample can then be produced by replacing the bracketed event role names with their corresponding value in the JSON object.
@@
We evaluate MLA on DocEE, which to the best of our knowledge is the only FSCD EAE dataset for long documents. Our experimental results show that MLA significantly improves performance on zero and few-shot roles while boosting overall F1 score. We additionally show that MLA improves performance on target domain roles which are semantically different from those observed in the source domain. To do so, we introduce a new metric, Role-Depth F1 (RDF1), which uses a statistical depth Seegmiller and Preum (2023) to identify roles in the target domain which are semantic outliers with respect to roles observed in the source domain. For example, the common source-domain role location has semantic similarity with other location-based roles, such as temporary settlement. Conversely, target roles such as maximum wind speed or Magnitude (Tsunami Heights) are domain specific and have little semantic overlap with the source training data. We visualize this phenomena in Figure 3. We argue that RDF1 provides a more useful assessment of cross-domain performance compared to classic F1, which does not capture such semantic nuances.
@@
    We introduce the first document-level EAE data augmentation framework, Mad Lib Aug (MLA). MLA can generate novel documents for DocEAE augmentation and is uniquely suited to facilitate generation of zero-shot, cross-domain event roles.
@@
    We show that MLA significantly improves performance on a diverse suite of few-shot cross-domain metrics, including a novel EAE evaluation metric, Role-Depth F1 (RDF1). RDF1 characterizes model performance on semantically outlying roles. We show that models trained using data generated with MLA yield significant improvement in RDF1, highlighting the value of MLA in cross-domain DocEAE task.
@@ 2.2 Data Augmentation for EAE
One concurrent work that uses LLMs to generate EAE data for sentence-level tasks is Ma et al. (2023), where they employ LLMs in a multi-step prompt and self-reflect pipeline. However, there are a number of reasons why we cannot apply this work as a baseline. First, their pipeline leverages information that is unavailable for the DocEE dataset, e.g., event definitions, role definitions, and valid entity types per role. Additionally, they provide the LLM with in-domain demonstrations for trigger generation, when our goal is to use no in-domain data for generation. Finally, their proposed pipeline uses significant LLM-based self-reflection. This is a concern as, for document-level tasks (i) self-reflection makes augmentation significantly more expensive when using OpenAI models and (ii) from our experience, self-reflection becomes less effective as document-length increases. Future works may explore how to adapt Ma et al. (2023) to document-level tasks.
@@ 2.3 Event Extraction Evaluation
Prior works on EAE in trigger-free datasets use evaluation metrics such as exact match F1 or head noun match F1 Tong et al. (2022). Recently, there have been some advancements in EAE evalulation for trigger-centric datasets. For example, Huang et al. (2023) proposed a new metric (ARG-I+) for trigger-based event extraction to better capture whether a role is associated with the correct trigger. In Li et al. (2022), they explicitly examine performance on long-tail role types. However, to the best of our knowledge, there are no related works on metrics which can measure an EAE model’s performance on semantically outlying roles during cross-domain evaluation. We propose a novel EAE evaluation metric, Role-Depth F1 (RDF1), to address this gap.
@@ 3 Methods
In this section we first describe Mad Lib Aug, our LLM-powered data augmentation framework for EAE document generation. Then, we formally introduce our novel metric, Role-Depth F1, which evaluates model performance on semantically outlying role types.
3.1 Mad Lib-based Augmentation
Mad Lib Aug (MLA) leverages the intuition that Mad Libs share a striking resemblance to EAE data. Mad Libs, first released in 1953, are a popular party activity in which a document has various masked phrases, each with a category such as “verb", “car brand", or “type of food". The goal of the game is to choose phrases for each category without knowing the context of the document. The final document is then read aloud, where the chosen category names often produce humorous stories.
The intuition regarding the connection between Mad Libs and EAE data is as follows. If a Mad Lib was written about an event, and used categories which were selected from that event ontology, then a context-aware solution to that Mad Lib would produce EAE annotation. This connection is powerful as LLMs such as GPT-3.5 have significant familiarity with Mad Libs and thus require little to no prompt optimization when requested to generate them. Additionally, Mad Libs are historically narrative, and thus easy to generate using varying styles. In this work, we break EAE data generation down into a two-step framework (i) Mad Lib Generation and (ii) Mad Lib Slot Filling. We describe each step in detail in the following subsections.
3.1.1 Mad Lib Generation
@@
Our Mad Lib generation framework can be found in the top-half of Figure 2. We prompt the LLM to write a Mad Lib about a given <event type> and list each role in the event’s ontology as a category. Thus for each sample, we aim to generate a document which contains every possible role for a given event. Our strategy does not generate documents which include/exclude certain event role subsets. However, we choose this strategy for the following two reasons. (i) In order to perform a fair comparison between augmented and real data in the FSCD setting, we only generate 5 documents per event type. Thus, given our goal is to improve performance on zero and few-shot roles, each document should be as role-dense as possible. (ii) A core challenge of document-level EAE data generation is the issue of generating under-labeled samples. That is, if the desired length of a document about a single event is large, but the number of roles to be included is small, it is very difficult to prevent the LLM from accidentally generating an unintended role. By providing the LLM with all possible target/desired roles as Mad Lib categories, we avoid this issue. To promote high-quality output format, our Mad Lib generation is done using $k=5$ in-context demonstrations, where samples are randomly sampled from a subset of the source domain documents. Specifically, we sample from role-dense documents of manageable length to manage API costs (see Appendix B.2 for details). These randomly sampled documents are then converted into Mad Lib format by replacing argument spans with role names. Note our method has the flexibility to promote document diversity by instructing the model to generate documents in a certain style, such as “newspaper article".
3.1.2 Mad Lib Slot Filling
Once we have generated a Mad Lib, we feed it back into the LLM as shown in the bottom-half of Figure 2. The LLM is instructed to generate a JSON object where the keys are the “categories", or event roles, and the values are the arguments. Note that we choose this format as OpenAI LLMs are able to generate JSON data  [^2] https://platform.openai.com/docs/api-reference/chat/create. We use the $k=5$ documents from the source domain as example data for the LLM in few-shot generation of the JSON object.
We finally do string replacement between the bracketed (category) names in the Mad Lib, and the corresponding key-value pair in the JSON object. The result is an event-centric document with annotation for argument extraction. Note that if during generation the model deviates from generating the exact category name (e.g. generating [the location] instead of [location]), we use SBERT Reimers and Gurevych (2019) to encode the mismatched category name and compute the cosine similarity between it and all possible role names for a given sample. We can reliably match the generated category to the true category using this method and then produce EAE annotation.
More details on Mad Lib Generation and Mad Lib Slot Filling can be found in Appendix B.2. See Appendix C for example augmented data.
@@ 3.2 Role-Depth-F1 (RDF1)
We embed each role in each set using S-BERT Reimers and Gurevych (2019)333sentence-transformers/all-MiniLM-L6-v2 to get a set of source role embeddings $S$ and a set of target role embeddings $T$. TTE depth scores each target role embedding $t\in T$ according to
@@
We consider the 25% of target domain roles with the lowest TTE depth scores to be "outlying" roles with respect to the source domain roles. By computing F1 only on these roles, i.e. RDF1, we measure the ability of a FSCD EAE model to perform domain transfer to roles which deviate semantically from the source domain.
In Figure 3, we visualize how depth scores rank all roles in the target domain with respect to the source domain. Notably, we see that Maximum Wind Speed, Damaged Crops & Livestock, and Magnitude (Tsunami heights) are the most semantically outlying, with influence people and temporary settlement being examples of roles closer to source domain roles. This plot justifies the need for RDF1 as a cross-domain evaluation metric, as it follows ones intuition that, for example, temporary settlement may have semantic overlap with source domain concepts like location. Note that this method will inherently filter out evaluation of any overlapping roles between source and target domain.
In DocEE, 13 roles comprise the set of semantic outliers evaluated in RDF1. 12/13 of these roles are few-shot (i.e. < 5 training examples), making RDF1 a more challenging version of the few-shot evaluation subset. We additionally note that the embedding strategy for creating $S$ and $T$ from $R_{S}$ and $R_{T}$ is flexible; event ontologies which include additional information such as role descriptions or valid entity types may be used to enrich the role representations. Since DocEE provides no such information, opting instead for use of semantically-rich role names, the role names themselves are used as $R_{S}$ and $R_{T}$.
4 Experiments
In this section, we describe the details of our experimental setup. Specifically, we introduce the (i) Dataset (ii) Baseline Models, (iii) Baseline Augmentation Methods, and (iv) Evaluation Metrics.
4.1 Dataset
We evaluate our method on the cross-domain split of the DocEE dataset. The DocEE dataset distribution is shown in Table 1. The task is structured to allow pre-training on a large number of source domain events. Then, the goal is to fine-tune this model to a target domain for which there are only very few annotated documents. In this dataset, the target domain is comprised of 10 event types, which are all under the common theme of “Natural Disasters". This grouping of source and target event domains was chosen to reduce overlap between source and target role types Tong et al. (2022). The target and source domains have a small 15% overlap, with common roles such as date, location, and economic loss occurring in both domains. During few-shot cross-domain training, only 5 documents per-event type are provided.
@@ 4.1 Dataset
Table 1: DocEE data distribution. The source domain contains annotation for over 23k documents across 49 different event types, with over 150k argument instances. The target domain contains annotation for documents with 10 distinct event types, all under the common theme of “Natural Disasters".
4.2 Baseline EAE Models
In this study, we evaluate the impact of EAE data augmentation using the following two EAE models adapted from the DocEE Paper Tong et al. (2022).
  • LongFormer-Seq: This model treats EAE as a token labeling task. Specifically, for a document $D$ with tokens $\{t_{0},\dots,t_{n}\}\in D$, we label argument spans using the label space $\{r_{0},\dots,r_{n}\}\in R$, where a given $r_{i}$ denotes an event role. Given that we are processing long documents, we use the LongFormer-base (149M parameters) Beltagy et al. (2020) architecture, which can process up to 4096 tokens. We use this model as it has been shown to out-perform BERT-based token classifiers on DocEAE Tong et al. (2022).
  • BERT-QA: This model treats EAE as an extractive question-answering task Du and Cardie (2020). In this formulation, the question is the argument role, and the context is the document. The goal is to predict the span(s) $(start_{i},end_{i})$ corresponding to the argument of the role in the document. Note that for this baseline, we use a BERT-base (110M parameters) Devlin et al. (2019) model pre-trained on the SQUAD 2.0 Rajpurkar et al. (2016) question-answering dataset. We choose this baseline as QA-based methods are common in EAE Du and Cardie (2020); Lu et al. (2023) and have been shown to perform well on DocEE in prior work Tong et al. (2022).

### v2

# Large Language Models for Document-Level Event-Argument Data Augmentation for Challenging Role Types

@@ Abstract
Event Argument Extraction (EAE) is an extremely difficult information extraction problem — with significant limitations in few-shot cross-domain (FSCD) settings. A common solution to FSCD modeling is data augmentation. Unfortunately, existing augmentation methods are not well-suited to a variety of real-world EAE contexts including (i) The need to model long documents (10+ sentences) (ii) The need to model zero and few-shot roles (i.e. event roles with little to no training representation). In this work, we introduce two novel LLM-powered data augmentation frameworks for synthesizing extractive document-level EAE samples using zero in-domain training data. Our highest performing methods provide a 16-pt increase in F1 score on extraction of zero shot role types.
To better facilitate analysis of cross-domain EAE, we additionally introduce a new metric, Role-Depth F1 (RDF1), which uses statistical depth to identify roles in the target domain which are semantic outliers with respect to roles observed in the source domain. Our experiments show that LLM-based augmentation can boost RDF1 performance by up to 11 F1 points compared to baseline methods.
Large Language Models for Document-Level Event-Argument Data Augmentation for Challenging Role Types
Joseph Gatto, Parker Seemiller, Omar Sharif, Sarah M. Preum Department of Computer Science, Dartmouth College joseph.m.gatto.gr@dartmouth.edu
@@ 1 Introduction
In recent years, there has been considerable progress in the domain of Event Argument Extraction (EAE), as advances in question answering Du and Cardie (2020), prompt tuning Ma et al. (2022), and semantic graph modeling Yang et al. (2023b) have led to state-of-the-art results on EAE benchmarks. However, some simplifying assumptions made in prior work limit their applicability in a real-world setting, including use of sentence-level data, reliance on large-scale annotation, or that training and testing data come from the same domain and thus share a similar label distribution.
@@
Figure 1: In (a) we leverage LLM’s prior knowledge of Mad Libs to generate templated, categorized documents. MadLib solutions can then be used as EAE annotation. (b) We generate event structure data for an event and use LLMs to convert the structure into an EAE document. We employ semantic n-gram matching to align the structure with spans in the generated document.
@@
A key issue in the FSCD DocEAE problem is the need to model 0-shot roles Yang et al. (2023a), i.e. roles which exist in an event’s schema and thus might appear in the test set, but have no representation in the training data. To illustrate this, consider the FSCD DocEE dataset Tong et al. (2022), where the task is to transfer knowledge from a collection of source domain events (e.g., Road Crash, Celebrity Death, Bank Robbery) to a set of events in a new domain, Natural Disasters (e.g., Droughts, Earthquakes, Tsunamis). Each natural disaster in the FSCD training set only has 5 documents available for training. Across all samples representing the Volcano Eruption event, for example, zero documents contain annotation for the role The State of the Volcano (Dormant or Active). This makes this role extremely challenging for many DocEAE models to extract.
@@
One solution to this problem is to strategically generate training samples in order to increase representations of zero and few-shot roles. However, to the best of our knowledge, no prior work has generated DocEAE training samples in a way that can help resolve the aforementioned issues. For example, many related works in EAE data augmentation such as Gao et al. (2022); Ma et al. (2023); Wang et al. (2023) only focus on sentence-level tasks. The few prior works in DocEAE augmentation do not generate new data samples but rather augment existing samples Liu et al. (2022) or use a pre-trained model to weakly label unannotated corpora Liu et al. (2021).
@@
In this paper, we address this issue by introducing two strategies for LLM-powered 0-shot DocEAE data augmentation: (i) Mad Lib Aug (MLA) is inspired by the insight that EAE samples can be formulated as the popular game Mad Libs, where the “blanks" found in Mad Libs correspond to roles in an event ontology. This allows us to leverage LLMs strong conceptual priors of Mad Libs for templated text generation. We visualize the relationship between Mad Libs and EAE in Figure 1. We use LLMs to both generate and solve Mad Libs in this framework. (ii) Struct2Text (S2T): is modeled off the synthetic data generation module provided by the LangChain library  [^1] https://www.langchain.com/, which can transform structured information (e.g. a dictionary of role-argument mappings) into documents. Note that both MLA and S2T require no in-domain training samples and can generate data for 0-shot roles given only the event title and role name thus facilitating generalizable cross-domain EAE.
@@
We evaluate each method on DocEE, which to the best of our knowledge is the only FSCD EAE dataset for long documents. Our experimental results show that each approach significantly improves performance on zero-shot roles without sacrificing overall F1 score. We additionally show performance gains on target domain roles which are semantically different from those observed in the source domain. To do so, we introduce a new metric, Role-Depth F1 (RDF1), which uses a statistical depth Seegmiller and Preum (2023) to identify roles in the target domain which are semantic outliers with respect to roles observed in the source domain. For example, the common source-domain role location has semantic similarity with other location-based roles, such as temporary settlement. Conversely, target roles such as maximum wind speed or Magnitude (Tsunami Heights) are domain specific and have little semantic overlap with the source training data. We visualize this phenomena in Figure 2. We argue that RDF1 provides a more useful assessment of cross-domain performance compared to classic F1, which does not capture such semantic nuances.
@@
    We establish two LLM-powered document-level EAE data augmentation strategies. Our methods can generate novel documents for DocEAE augmentation and are uniquely suited to facilitate generation of zero-shot, cross-domain event roles. Our approach is generalizable to future low-resource event data, as only the only in-domain information needed for document generation is the event schema.
@@
    We show that DocEAE augmentation significantly improves performance on a comprehensive suite of EAE metrics, including a novel evaluation metric, Role-Depth F1 (RDF1). RDF1 characterizes model performance on semantically outlying roles. We show that models trained using data generated with MLA and S2T yield significant improvement in RDF1, highlighting the value of augmentation in the cross-domain DocEAE task — paving the way for future work in this space.
@@ 2.2 Data Augmentation for EAE
One concurrent work that uses LLMs to generate EAE data for sentence-level tasks is Ma et al. (2023), where they employ LLMs in a multi-step prompt and self-reflect pipeline. However, there are a number of reasons why we cannot apply this work as a baseline. First, their pipeline leverages information that is unavailable for the DocEE dataset, e.g., event definitions, role definitions, and valid entity types per role. Additionally, they provide the LLM with in-domain demonstrations for trigger generation, when our goal is to use no in-domain data for generation. Finally, their proposed pipeline uses significant LLM-based self-reflection. This is a concern as, for document-level tasks (i) self-reflection makes augmentation significantly more expensive when using proprietary LLMs and (ii) from our experience, self-reflection becomes less effective as document-length increases. Future works may explore how to adapt Ma et al. (2023) to document-level tasks.
@@ 2.3 Event Extraction Evaluation
Prior works on EAE in trigger-free datasets use evaluation metrics such as exact match F1 or head noun match F1 Tong et al. (2022). Recently, there have been some advancements in EAE evalulation for trigger-centric datasets. For example, Huang et al. (2023) proposed a new metric (ARG-I+) for trigger-based event extraction to better capture whether a role is associated with the correct trigger. In Li et al. (2022), they explicitly examine performance on long-tail role types. However, to the best of our knowledge, there are no related works on metrics which can measure an EAE model’s performance on semantically outlying roles during cross-domain evaluation. It is crucial to understand to what degree a cross-domain learning algorithm can model truly novel events and not just those which are semantically similar to prior seen events. We propose a novel EAE evaluation metric, Role-Depth F1 (RDF1), to address this gap.
@@ 3 Methods
In this section we describe both Mad Lib Aug and Struct2Text. Then, we formally introduce our novel metric, Role-Depth F1, which evaluates model performance on semantically outlying role types.
3.1 Mad Lib Aug (MLA)
Mad Lib Generation and Slot-Filling:
@@ 3.1 Mad Lib Aug (MLA)
Recall that Mad Libs are texts interwoven with categorized missing information slots (See Figure 1 for an example). Thus, to generate an event-driven Mad Lib, we prompt the LLM to write a Mad Lib about a given event type and list each role in the event’s schema as a category. Thus for each sample, we aim to generate a document which contains every possible role for a given event. Our strategy does not generate documents which intentionally include/exclude certain event role subsets. We choose this strategy for the following reason: A core challenge of document-level EAE data generation is the issue of generating under-labeled samples. That is, if the desired length of a document about a single event is large, but the number of roles to be included is small, it is very difficult to prevent the LLM from accidentally generating an unintended role. By providing the LLM with all possible target/desired roles as Mad Lib categories, we mitigate this issue.
Once we have generated a Mad Lib, we ask the LLM to fill-in the categories (slot-filling) with appropriate information given the document. This is done by prompting the LLM to return a list of role : argument pairs which can be slotted into the templated document. We finally do string replacement between the bracketed category names in the Mad Lib, and the LLM-generated solution (shown in Figure 1). The result is an event-centric document with annotation for argument extraction. In the event where the LLM generates a Mad Lib with imperfections such as large number of missing arguments or hallucinated roles, we perform a post processing step which is discussed in detail in Appendix B.3.
For both Mad Lib Generation and Slot-Filling, we use k=3 documents from the source domain to help the model understand the input/output format. We choose k=3 samples with moderate length and high role-density to minimize API costs for use with proprietary models. Generated data statistics can be found in Table LABEL:gen_data_stats, where we show MLA produces documents with sentence length between 8.2-12.2 across four different LLM backbones.
@@
3.2 Struct2Text (S2T)
This approach modifies the data augmentation tool provided by LangChain which aims to convert structured data into text. At a high-level, this method uses OpenAI’s function calling capability to populate a Python class with event information. We thus create a custom class for each event in the target domain, where the class name is the event name and each role in the event is a class member of type string. We closely follow the synthetic data generation strategy implemented by LangChain  [^3] https://python.langchain.com/docs/use_cases/data_generation/ with a few key modifications (i) We alter the prompt structure to encourage writing of document-level data (i.e. 10-sentence long documents). (ii) We instruct the model to write the documents in the style of a newspaper article to help match the DocEE target domain (iii) We perform semantic n-gram matching to align event structures with generated documents. This is a necessary step as Struct2Text, when used off-the-shelf, does not reliably produce documents with the exact substrings requested when producing document-level data. Our matching algorithm compares the argument in the synthetic event structure with every possible n-gram (up to n=20) in the document and returns the n-gram with the highest cosine similarity using SBERT 444sentence-transformers/all-MiniLM-L6-v2. We find that for semantic similarity greater than 0.5, this method reliably aligns structure to text. Any arguments which have cosine similarity lower than our threshold are considered to have not been generated and are discarded from the synthetic event record. See Table LABEL:gen_data_stats for S2T generated data statistics.
@@ 3.3 Role-Depth-F1 (RDF1)
We embed each role in each set using S-BERT Reimers and Gurevych (2019) to get a set of source role embeddings $S$ and a set of target role embeddings $T$. TTE depth scores each target role embedding $t\in T$ according to
@@
We consider the 25% of target domain roles with the lowest TTE depth scores to be "outlying" roles with respect to the source domain roles. We choose 25% for this TTE depth threshold as it strikes a good balance between filtering for target roles which differ from source domain roles, while maintaining a good amount of roles for evaluation. We explore how the choice of RDF1 threshold affects performance in Table 3 in the Appendix. By computing F1 only on these roles, i.e. RDF1, we measure the ability of a FSCD EAE model to perform domain transfer to roles which deviate semantically from the source domain.
In Figure 2, we visualize how depth scores rank all roles in the target domain with respect to the source domain. Notably, we see that Maximum Wind Speed, Damaged Crops & Livestock, and Magnitude (Tsunami heights) are the most semantically outlying, with influence people and temporary settlement being examples of roles closer to source domain roles. This plot justifies the need for RDF1 as a cross-domain evaluation metric, as it follows ones intuition that, for example, temporary settlement may have semantic overlap with source domain concepts like location.
In DocEE, 13 roles comprise the set of semantic outliers evaluated in RDF1. 12/13 of these roles have less than 5 training examples and can be considered few-shot roles. We additionally note that the embedding strategy for creating $S$ and $T$ from $R_{S}$ and $R_{T}$ is flexible; event ontologies which include additional information such as role descriptions or valid entity types may be used to enrich the role representations. Since DocEE provides no such information, opting instead for use of semantically-rich role names, the role names themselves are used as $R_{S}$ and $R_{T}$.
@@
Table 1: DocEE data distribution. The source domain contains annotation for over 23k documents across 49 different event types. The target domain contains annotation for documents with 10 distinct event types, all under the common theme of “Natural Disasters".
4 Experiments
In this section, we describe the details of our experimental setup. Specifically, we introduce the (i) Dataset (ii) EAE Models, (iii) Baseline Augmentation Strategies, and (iv) Evaluation Metrics.
4.1 Dataset
We evaluate our method on the cross-domain split of the DocEE dataset. The DocEE dataset distribution is shown in Table 1. The task is structured to allow pre-training on a large number of source domain events. The source domain in DocEE covers a wide variety of events including Famous Person - Death, Resgination / Dismissal, and Election, amongst others. The goal is to fine-tune this model to a target domain for which there are only very few annotated documents. In this dataset, the target domain is comprised of 10 event types, which are all under the common theme of “Natural Disasters". This grouping of source and target event domains was chosen to reduce overlap between source and target role types Tong et al. (2022). The target and source domains have a small 15% overlap, with common roles such as date, location, and economic loss occurring in both domains. During few-shot cross-domain training, only 5 documents per-event type are provided.
4.2 EAE Models
In this study, we evaluate the impact of EAE data augmentation using the following two EAE models motivated by the DocEE Paper Tong et al. (2022).
  • LongFormer(LF)-Seq: This model treats EAE as a token labeling task. Specifically, for a document $D$ with tokens $\{t_{0},\dots,t_{n}\}\in D$, we label argument spans using the label space $\{r_{0},\dots,r_{n}\}\in R$, where a given $r_{i}$ denotes an event role. Given that we are processing long documents, we use the LongFormer-base (149M parameters) Beltagy et al. (2020) architecture, which can process up to 4096 tokens. We use this model as it has been shown to out-perform BERT-based token classifiers on DocEAE Tong et al. (2022).
  • BERT-QA: This model treats EAE as an extractive question-answering task, inspired by Du and Cardie (2020). In this formulation, the question is the argument role, and the context is the document. The goal is to predict the span(s) $(start_{i},end_{i})$ corresponding to the argument of the role in the document. Note that for this baseline, we use a BERT-base (110M parameters) Devlin et al. (2019) model as in the original DocEE paper.